# Semilinear spaces of matrices with transitivity, rank or spectral constraints

### v1

@@ 1.4 Main new results on trivial spectrum subspaces
Let $\mathbb{D}$ be a division ring with degree $d$ over a field $\mathbb{F}$, and $V$ be a (right) vector space over $\mathbb{D}$ with finite dimension $n>0$. Assume that $|\mathbb{F}|\geq dn$. Then the greatest possible dimension for an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ (or, equivalently, of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$) is
$\alpha(n,d)=d\,\dbinom{n}{2}+n(d-1).$
@@
An immediate consequence is that when $|\mathbb{F}|\geq dn$ the composition problem stated in the end of Section 1.2 has a positive answer in $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$: Indeed in that case $|\mathbb{F}|\geq dk$ for all $k\in\mathopen{[\![}1,n\mathclose{]\!]}$, and for every partition $n=n_{1}+\cdots+n_{p}$ the number $\sum_{k=1}^{p}\alpha(n_{k},d)+d\sum_{1\leq i<j\leq p}n_{i}n_{j}$ equals $\alpha(n,d)$.
Under the same assumption as in Theorem 1.6, the classification of the irreducible optimal spaces is partly given by the following theorem:
@@
Theorem 1.7 (Partial classification theorem for irreducible optimal $\mathbb{F}$-trivial spectrum subspaces).
@@
    If $n=1$ then the irreducible optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ are the $\mathbb{F}$-linear hyperplanes of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that do not contain $\operatorname{\operatorname{id}}_{V}$.
@@
Of course, the above two theorems look on the surface as plain generalizations of the case $\mathbb{D}=\mathbb{F}$, and in some sense they are. But there is a subtlety: although $\mathcal{A}_{b,\mathbb{D}}$ is always an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, in general it is not an optimal one! Hence it is necessary to understand when optimality happens. Some additional terminology is necessary for this.
@@
In cases (i) and (ii), we denote by $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$ the standard trace of $\mathbb{D}$ over $\mathbb{F}$ (see Section 2.1 for details), and in case (iii) we denote by $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$ the reduced trace $x\mapsto x+\sigma(x)$ (the reduced trace of $x$ is either $2x$ if $x$ belongs to $\mathbb{F}1_{\mathbb{D}}$, otherwise it equals the opposite of the coefficient on $t$ of the minimal polynomial of $x$ over $\mathbb{F}$).
@@
Finally, if $\sigma$ is an antiautomorphism of the $\mathbb{F}$-algebra $\mathbb{D}$, we define a sesquilinear form on $V$ as a mapping $B:V^{2}\rightarrow\mathbb{D}$ that is right-linear and left-$\sigma$-quasilinear, meaning that $B(\alpha x+y,-)=\sigma(\alpha)B(x,-)+B(y,-)$ and $B(-,\alpha x+y)=\alpha B(-,x)+B(-,y)$ for all $\alpha\in\mathbb{D}$ and $x,y$ in $V$. Two such forms $B$ and $B^{\prime}$ are called equivalent if there exists an automorphism $\varphi$ of the $\mathbb{D}$-vector space $V$ such that $B^{\prime}(\varphi(x),\varphi(y))=B(x,y)$ for all $(x,y)\in V^{2}$.
Theorem 1.8 (Full classification theorem for optimal $\mathbb{F}$-trivial spectrum subspaces).
Let $V$ be a (right) $\mathbb{D}$-vector space with finite dimension $n>0$. Assume that $|\mathbb{F}|\geq nd$, and let $b$ be a nonisotropic $\mathbb{F}$-bilinear form on $V$. For $\mathcal{A}_{b,\mathbb{D}}$ to be an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, it is necessary and sufficient that all the following conditions hold:
@@
    Given an associated pair $(\sigma,e)$ for $\mathbb{D}$ over $\mathbb{F}$, there is a $\sigma$-sesquilinear form $B$ on the $\mathbb{D}$-vector space $V$ such that
    $\forall(x,y)\in V^{2},\;b(x,y)=e(B(x,y)).$
@@
Hence, if $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ with associated pair $(\sigma,e)$, and we take any $\sigma$-sesquilinear form $B$ on $V$, the space $\mathcal{A}_{b,\mathbb{D}}$ is an irreducible $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ with dimension $\alpha(n,d)$ provided that $B$ is $e$-nonisotropic , meaning that $\forall x\in V,\;e(B(x,x))\neq 0$ (for this very result the provision $|\mathbb{F}|\geq nd$ is not needed, see the proof in Section 5.6).
As a consequence of Theorems 1.7 and 1.8, and still under the cardinality assumption from Theorem 1.6, we can refine our understanding of the irreducible optimal trivial spectrum subspaces:
@@
Assume that $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$. Let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, with $|\mathbb{F}|\geq nd$. The space $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ contains an irreducible optimal $\mathbb{F}$-trivial spectrum subspace only if $n=1$.
@@
The next result deals with the conjugacy class of $\mathcal{A}_{b,\mathbb{D}}$ inside the algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
@@
Let $V$ be a (right) $\mathbb{D}$-vector space with finite dimension $n>0$. Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, with associated pair $(\sigma,e)$. Let $B$ and $B^{\prime}$ be $e$-nonisotropic $\sigma$-sesquilinear forms on $V$, and set $b:(x,y)\mapsto e(B(x,y))$ and $b^{\prime}:(x,y)\mapsto e(B^{\prime}(x,y))$. Then $\mathcal{A}_{b,\mathbb{D}}$ is conjugated to $\mathcal{A}_{b^{\prime},\mathbb{D}}$ in the algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ if and only if $B^{\prime}$ is equivalent to $\alpha B$ as a $\sigma$-sesquilinear form for some $\alpha\in\mathbb{F}^{\times}$.
@@
Conversely, let $M\in\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$. By the same computation as in the above, we see that the identity $M^{\star}=-M$ leads to $X\in(\mathbb{D}^{n})\mapsto e(X^{\star}MX)$ being additive (its polar form vanishes). As a consequence, for all $X=\begin{bmatrix}x_{1}&\cdots&x_{n}\end{bmatrix}^{T}\in\mathbb{D}^{n}$, we find
@@
    The irreducible optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are the spaces of the form $P^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ where $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ is $e$-nonisotropic.
@@
    Given $e$-nonisotropic matrices $P_{1}$ and $P_{2}$ in $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$, the spaces $P_{1}^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ and $P_{2}^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ are conjugated under an element of $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ if and only if $P_{2}$ is star-congruent to $\alpha P_{1}$ for some $\alpha\in\mathbb{F}^{\times}$ (two matrices $M$ and $M^{\prime}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are star-congruent if and only if there exists $Q\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ such that $M^{\prime}=QMQ^{\star}$).
@@
Classically $\mathbb{D}:\mathbb{F}]=2^{p}$, $\mathbb{F}:\mathbb{K}]=2^{n}$ and $\mathbb{K}$ is the set of all squares of elements of $\mathbb{D}$; moreover $X_{1},\dots,X_{n}$ are linearly independent over $\mathbb{K}$. Now, take an arbitrary $\mathbb{F}$-linear form $e$ on $\mathbb{D}$ that maps $X_{1},\dots,X_{n}$ to themselves, and consider the diagonal matrix $P:=\operatorname{\operatorname{Diag}}(X_{1},\dots,X_{n})$. For all $Y=\begin{bmatrix}y_{1}&\cdots&y_{n}\end{bmatrix}^{T}\in\mathbb{D}^{n}$ we have $e(Y^{T}PY)=\sum_{i=1}^{p}(y_{i})^{2}X_{i}$, which vanishes if and only if $y_{1}=\cdots=y_{n}=0$. Hence $P$ is $e$-nonisotropic. Therefore $P^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ is an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, and both $n$ and $p$ can be taken as large as we wish.
Let us conclude by combining Theorem 1.2 with the previous results, so as to give a full description of the optimal $\mathbb{F}$-trivial spectrum spaces. We distinguish between two cases, whether $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ or not.
@@
Assume that $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$. Let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, with $|\mathbb{F}|\geq nd$. Let $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$. Then the $\mathcal{S}$-invariant $\mathbb{D}$-linear subspaces form a complete flag $(V_{0},\dots,V_{n})$ of $V$, and $\mathcal{S}=H_{1}\vee\cdots\vee H_{n}$, where for each $i\in\mathopen{[\![}1,n\mathclose{]\!]}$, $H_{i}$ is an $\mathbb{F}$-linear hyperplane of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{i}/V_{i-1})$ that does not contain the identity of $V_{i}/V_{i-1}$.
Conversely, given a complete flag $(V_{0},\dots,V_{n})$ of $V$, and for all $i\in\mathopen{[\![}1,n\mathclose{]\!]}$ an $\mathbb{F}$-linear hyperplane $H_{i}$ of $V_{i}/V_{i-1}$ that does not contain the identity of $V_{i}/V_{i-1}$, the space $H_{1}\vee\cdots\vee H_{n}$ is an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
@@
The optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are the spaces that are conjugated to a space of the form
$H_{1}\vee\cdots\vee H_{n}=\left\{\begin{bmatrix}\alpha_{1}&&[*]\\ &\ddots&\\ [0]&&\alpha_{n}\end{bmatrix}\mid(\alpha_{1},\dots,\alpha_{n})\in H_{1}\times\cdots\times H_{n}\right\}$
@@
where $H_{1},\dots,H_{n}$ are $\mathbb{F}$-linear hyperplanes of $\mathbb{D}$ that do not contain $1_{\mathbb{D}}$. Two such spaces $H_{1}\vee\cdots\vee H_{n}$ and $H^{\prime}_{1}\vee\cdots\vee H^{\prime}_{n}$ are conjugated in the algebra $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ if and only if, for all $i\in\mathopen{[\![}1,n\mathclose{]\!]}$, the hyperplanes $H_{i}$ and $H^{\prime}_{i}$ are conjugated in $\mathbb{D}$.
Theorem 1.16 (Full classification theorem for optimal $\mathbb{F}$-trivial spectrum subspaces).
@@
Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, with associated pair $(\sigma,e)$. Let $V$ be a $\mathbb{D}$-vector space with finite dimension $n>0$, where $|\mathbb{F}|\geq nd$, and $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
@@
Then the $\mathcal{S}$-invariant $\mathbb{D}$-linear subspaces constitute a (potentially incomplete) flag $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ there exists an $e$-nonisotropic $\sigma$-sesquilinear form $B_{i}$ on $V_{i}/V_{i-1}$ such that, for $b_{i}:=e\circ B_{i}$,
$\mathcal{S}=\mathcal{A}_{b_{1},\mathbb{D}}\vee\cdots\vee\mathcal{A}_{b_{p},\mathbb{D}}.$
@@
Moreover, each form $B_{i}$ is uniquely determined by $\mathcal{S}$ up to multiplication by an element of $\mathbb{F}^{\times}$, and the equivalence class of $B_{i}$ is uniquely determined by the conjugacy class of $\mathcal{S}$ in $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ up to multiplication by an element of $\mathbb{F}^{\times}$.
@@
Conversely, given a (potentially incomplete) flag $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ an $e$-nonisotropic $\sigma$-sesquilinear form $B_{i}$ on $V_{i}/V_{i-1}$, with associated $\mathbb{F}$-bilinear form $b_{i}:=e\circ B_{i}$, the space $\mathcal{A}_{b_{1},\mathbb{D}}\vee\cdots\vee\mathcal{A}_{b_{p},\mathbb{D}}$ is an optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
@@
Theorem 1.17 (Classification theorem for optimal $\mathbb{F}$-trivial spectrum subspaces).
@@
Let $n\geq 1$ be such that $|\mathbb{F}|\geq nd$. For every optimal $\mathbb{F}$-trivial spectrum subspace $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, there exists a partition $n=n_{1}+\cdots+n_{p}$ into positive integers and a list $(P_{1},\dots,P_{p})\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{D})\times\cdots\times\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{D})$ of $e$-nonisotropic matrices such that

### v2

@@ 1.4 The new results on trivial spectrum subspaces
Let $\mathbb{D}$ be a division ring with dimension $d$ over a central subfield $\mathbb{F}$, and $V$ be a (right) vector space over $\mathbb{D}$ with finite dimension $n>0$. Assume that $|\mathbb{F}|\geq dn$. Then the greatest possible dimension for a trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ (or, equivalently, of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$) is
$\alpha(n,d)=n(d-1)+d\,\dbinom{n}{2}$.
@@
An immediate consequence is that when $|\mathbb{F}|\geq dn$ the reconstruction problem stated in the end of Section 1.2 has a positive answer in $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$: Indeed in that case $|\mathbb{F}|\geq dk$ for all $k\in\mathopen{[\![}1,n\mathclose{]\!]}$, and for every partition $n=n_{1}+\cdots+n_{p}$ the number $\sum_{k=1}^{p}\alpha(n_{k},d)+d\sum_{1\leq i<j\leq p}n_{i}n_{j}$ equals $\alpha(n,d)$.
Next, we turn to the structure of the optimal spaces. Here, we consider an $\mathbb{F}$-bilinear form $b:V^{2}\rightarrow\mathbb{F}$ for a $\mathbb{D}$-vector space $V$, and set
$\mathcal{A}_{b,\mathbb{D}}:=\mathcal{A}_{b}\cap\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$
(which is much smaller than the full space $\mathcal{A}_{b}$ as we are only considering $\mathbb{D}$-linear endomorphisms).
@@
Theorem 1.9 (Partial classification theorem for irreducible optimal trivial spectrum subspaces).
@@
    If $n=1$ then the irreducible optimal trivial spectrum subspaces of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ are the $\mathbb{F}$-linear hyperplanes of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that do not contain $\operatorname{\operatorname{id}}_{V}$.
@@
Of course, the above two theorems look on the surface as plain generalizations of the case $\mathbb{D}=\mathbb{F}$, and in some sense they are. But there is a subtlety: although $\mathcal{A}_{b,\mathbb{D}}$ is always a trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, in general it is not an optimal one! Hence it is necessary to understand when optimality happens. Some additional terminology is necessary for this.
@@
In cases (i) and (ii), we denote by $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$ the Galois trace of $\mathbb{D}$ over $\mathbb{F}$, and in case (iii) we denote by $\operatorname{\operatorname{tr}}_{\mathbb{D}/\mathbb{F}}$ the reduced trace $x\mapsto x+\sigma(x)$ (the reduced trace of $x$ is either $2x$ if $x$ belongs to $\mathbb{F}1_{\mathbb{D}}$, otherwise it equals the opposite of the coefficient on $t$ of the minimal polynomial of $x$ over $\mathbb{F}$).
@@
Finally, if $\sigma$ is an antiautomorphism of the $\mathbb{F}$-algebra $\mathbb{D}$, we define a $\sigma$-sesquilinear on $V$ as a mapping $B:V^{2}\rightarrow\mathbb{D}$ that is right-linear and left-$\sigma$-quasilinear, meaning that $B(x\alpha+y,-)=\sigma(\alpha)B(x,-)+B(y,-)$ and $B(-,x\alpha+y)=B(-,x)\alpha+B(-,y)$ for all $\alpha\in\mathbb{D}$ and $x,y$ in $V$. Two such forms $B$ and $B^{\prime}$ are called equivalent if there exists an automorphism $\varphi$ of the $\mathbb{D}$-vector space $V$ such that $B^{\prime}(\varphi(x),\varphi(y))=B(x,y)$ for all $(x,y)\in V^{2}$.
Theorem 1.10 (Full classification theorem for optimal trivial spectrum subspaces).
Let $V$ be a (right) $\mathbb{D}$-vector space with finite dimension $n>0$. Assume that $|\mathbb{F}|\geq nd$, and let $b$ be a nonisotropic $\mathbb{F}$-bilinear form on $V$. For $\mathcal{A}_{b,\mathbb{D}}$ to be an irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, it is necessary and sufficient that both of the following conditions hold:
@@
    Given an associated pair $(\sigma,e)$ for $\mathbb{D}$ over $\mathbb{F}$, there is a $\sigma$-sesquilinear form $B$ on the $\mathbb{D}$-vector space $V$ such that (x,y)V^2,   b(x,y)=e(B(x,y)).
@@
Hence, if $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ with associated pair $(\sigma,e)$, and we take any $\sigma$-sesquilinear form $B$ on $V$, the space $\mathcal{A}_{b,\mathbb{D}}$ is an irreducible trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ with dimension $\alpha(n,d)$ provided that $B$ is $e$-nonisotropic, meaning that $\forall x\in V,\;e(B(x,x))\neq 0$ (for this very result the provision $|\mathbb{F}|\geq nd$ is not needed, see the proof in Section 5.6).
As a consequence of Theorems 1.9 and 1.10, and still under the cardinality assumption from Theorem 1.8, we can refine our understanding of the irreducible optimal trivial spectrum subspaces:
@@
Assume that $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$. The space $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ contains an irreducible optimal trivial spectrum subspace only if $n=1$.
@@
The next result deals with the similarity class of $\mathcal{A}_{b,\mathbb{D}}$ inside the algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
@@
Let $V$ be a (right) $\mathbb{D}$-vector space with finite dimension $n>0$. Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, with associated pair $(\sigma,e)$. Let $B$ and $B^{\prime}$ be $e$-nonisotropic $\sigma$-sesquilinear forms on $V$, and set $b:(x,y)\mapsto e(B(x,y))$ and $b^{\prime}:(x,y)\mapsto e(B^{\prime}(x,y))$. Then $\mathcal{A}_{b,\mathbb{D}}$ is similar to $\mathcal{A}_{b^{\prime},\mathbb{D}}$ in the algebra $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ if and only if $B^{\prime}$ is equivalent to $\alpha B$ as a $\sigma$-sesquilinear form for some $\alpha\in\mathbb{F}^{\times}$.
@@
Conversely, let $M\in\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$. By the same computation as in the above, we see that the identity $M^{\star}=-M$ leads to $X\in\mathbb{D}^{n}\mapsto e(X^{\star}MX)$ being additive (its polar form vanishes). As a consequence, for all $X=\begin{bmatrix}x_{1}&\cdots&x_{n}\end{bmatrix}^{T}\in\mathbb{D}^{n}$, we find
@@
Finally, two matrices $A$ and $B$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ will be called homo-star-congruent whenever $B=\alpha\,PAP^{\star}$ for some $\alpha\in\mathbb{F}^{\times}$ and some $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$.
@@
    The irreducible optimal trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are the spaces of the form $P^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ where $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ is $e$-nonisotropic.
@@
    Given $e$-nonisotropic matrices $P_{1}$ and $P_{2}$ in $\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$, the spaces $P_{1}^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ and $P_{2}^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ are similar if and only if $P_{2}$ is homo-star-congruent to $P_{1}$.
@@
Classically $\mathbb{D}:\mathbb{F}]=2^{p}$, $\mathbb{F}:\mathbb{K}]=2^{n}$ and $\mathbb{K}$ is the set of all squares of elements of $\mathbb{D}$; moreover $X_{1},\dots,X_{n}$ are linearly independent over $\mathbb{K}$. Now, take an arbitrary $\mathbb{F}$-linear form $e$ on $\mathbb{D}$ that maps $X_{1},\dots,X_{n}$ to themselves, and consider the diagonal matrix $P:=\operatorname{\operatorname{Diag}}(X_{1},\dots,X_{n})$. For all $Y=\begin{bmatrix}y_{1}&\cdots&y_{n}\end{bmatrix}^{T}\in\mathbb{D}^{n}$ we have $e(Y^{T}PY)=\sum_{i=1}^{p}(y_{i})^{2}X_{i}$, which vanishes if and only if $y_{1}=\cdots=y_{n}=0$. Hence $P$ is $e$-nonisotropic. Therefore $P^{-1}\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$ is an irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, and both $n$ and $p$ can be taken as large as we wish.
Let us conclude by combining Theorem 1.1 with the previous results, so as to give a full description of the optimal trivial spectrum spaces. We distinguish between two cases, whether $\mathbb{D}$ is of quadratic type over $\mathbb{F}$ or not.
@@
Assume that $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$. Let $V$ be an $n$-dimensional vector space over $\mathbb{D}$, with $|\mathbb{F}|\geq nd$. Let $S$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$. Then the $S$-invariant $\mathbb{D}$-linear subspaces form a complete flag $(V_{0},\dots,V_{n})$ of $V$, and $S=H_{1}\vee\cdots\vee H_{n}$, where for each $i\in\mathopen{[\![}1,n\mathclose{]\!]}$, $H_{i}$ is an $\mathbb{F}$-linear hyperplane of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{i}/V_{i-1})$ that does not contain the identity of $V_{i}/V_{i-1}$.
Conversely, given a complete flag $(V_{0},\dots,V_{n})$ of $V$, and for all $i\in\mathopen{[\![}1,n\mathclose{]\!]}$ an $\mathbb{F}$-linear hyperplane $H_{i}$ of $V_{i}/V_{i-1}$ that does not contain the identity of $V_{i}/V_{i-1}$, the space $H_{1}\vee\cdots\vee H_{n}$ is an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
@@
The optimal trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are the spaces that are similar to a space of the form $H_{1}\vee\cdots\vee H_{n}$ where $H_{1},\dots,H_{n}$ are $\mathbb{F}$-linear hyperplanes of $\mathbb{D}$ that do not contain $1_{\mathbb{D}}$. Two such spaces $H_{1}\vee\cdots\vee H_{n}$ and $H^{\prime}_{1}\vee\cdots\vee H^{\prime}_{n}$ are similar in the algebra $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ if and only if, for all $i\in\mathopen{[\![}1,n\mathclose{]\!]}$, the hyperplanes $H_{i}$ and $H^{\prime}_{i}$ are conjugated in $\mathbb{D}$.
Hence, if $\mathbb{D}$ is not of quadratic type over $\mathbb{F}$ (and the cardinality of $\mathbb{F}$ is large enough) the classification of optimal trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ is reduced to the classification of the $\mathbb{F}$-linear hyperplanes of $\mathbb{D}$ that do not contain $1_{\mathbb{D}}$, up to conjugation. The latter classification is directly connected with the classification of conjugacy classes of elements in $\mathbb{D}$ by duality. To see this, consider the composite $\operatorname{\operatorname{Tr}}:\mathbb{D}\rightarrow\mathbb{F}$ of the reduced trace of $\mathbb{D}$ over its center $C$ and of the Galois trace $\operatorname{\operatorname{Tr}}_{C/\mathbb{F}}$: this is a nonzero $\mathbb{F}$-linear mapping that satisfies $\operatorname{\operatorname{Tr}}(xy)=\operatorname{\operatorname{Tr}}(yx)$ for all $x,y$ in $\mathbb{D}$. Then $a\mapsto[b\mapsto\operatorname{\operatorname{Tr}}(ab)]$ induces an $\mathbb{F}$-linear isomorphism from $\mathbb{D}$ to the dual space of the $\mathbb{F}$-vector space $\mathbb{D}$. It is then easily checked that this yields a bijective correspondence between the $\mathbb{F}$-linear hyperplanes of $\mathbb{D}$ and the equivalence classes in $\mathbb{D}^{\times}$ for the relation $a\cong b$ defined by $\exists\lambda\in\mathbb{F}^{\times}:\exists u\in\mathbb{D}^{\times}:b=\lambda\,uau^{-1}$. The $\mathbb{F}$-linear hyperplanes that do not contain $1_{\mathbb{D}}$ correspond to the elements of $\mathbb{D}$ with non-vanishing trace. Finally, by the Skolem-Noether theorem, two elements of $\mathbb{D}$ are conjugated if and only if they have the same minimal polynomial over the center $C$.
We finish with the case where $\mathbb{D}$ is of quadratic type over $\mathbb{F}$.
Theorem 1.18.
@@
Assume that $\mathbb{D}$ is of quadratic type over $\mathbb{F}$, with associated pair $(\sigma,e)$. Let $V$ be a $\mathbb{D}$-vector space with finite dimension $n>0$, where $|\mathbb{F}|\geq nd$, and $S$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
@@
Then the $S$-invariant $\mathbb{D}$-linear subspaces constitute a (potentially incomplete) flag $(V_{0},V_{1},\dots,V_{p})$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ there exists an $e$-nonisotropic $\sigma$-sesquilinear form $B_{i}$ on $V_{i}/V_{i-1}$ such that, for $b_{i}:=e\circ B_{i}$,
$S=\mathcal{A}_{b_{1},\mathbb{D}}\vee\cdots\vee\mathcal{A}_{b_{p},\mathbb{D}}.$
@@
Moreover, each form $B_{i}$ is uniquely determined by $S$ up to multiplication by an element of $\mathbb{F}^{\times}$, and the equivalence class of $B_{i}$ is uniquely determined by the similarity class of $S$ in $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ up to multiplication by an element of $\mathbb{F}^{\times}$.
@@
Conversely, given a (potentially incomplete) flag $(V_{0},V_{1},\dots,V_{p})$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ an $e$-nonisotropic $\sigma$-sesquilinear form $B_{i}$ on $V_{i}/V_{i-1}$, with associated $\mathbb{F}$-bilinear form $b_{i}:=e\circ B_{i}$, the space $\mathcal{A}_{b_{1},\mathbb{D}}\vee\cdots\vee\mathcal{A}_{b_{p},\mathbb{D}}$ is an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$.
@@
Theorem 1.19.
@@
Let $n\geq 1$ be such that $|\mathbb{F}|\geq nd$. For every optimal trivial spectrum subspace $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, there exists a partition $n=n_{1}+\cdots+n_{p}$ into positive integers and a list $(P_{1},\dots,P_{p})\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{D})\times\cdots\times\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{D})$ of $e$-nonisotropic matrices such that